OpenStudy (anonymous):

How many grams of carbon are present in 45.0 g of CCl4?

OpenStudy (anoop27):

CCl4=(12+35.5*4)gms=(12+142)gm=154gms 154 gm of CCl4 contains 12 gm of Carbon 1gm of CCl4 contains 12/154 gm of Carbon 45 gm of CCl4 contains (12*45)/154 gm of Carbon so your Answer is (12*45)/154 gms